OpenStudy (anonymous):

0.5(c+2.8)-c=6.8c+0.3

OpenStudy (anonymous):

0.5 c+1.4-c=6.8c+0.3 -0.5c+1.4=6.8c+0.3 +0.5c +0.5c-0.3 --------------------- 1.1=6.8c 1.1/6.8=6.8c/6.8 c=1.1/6.8